The table below gives the scores of Students on a English Examinations


under KCSE


Student


1


2


3


4


5


6


7


8


9


Oral test


B-


B+


B


A-


A


B


C+


B


C-


Composition


10


21


22


19


17


14


13


16


12


Literature


34


76


74


60


68


44


45


53


43


(Composition test is out of 30 and literature test is out of 100 marks)


(a) Calculate the value of the most appropriate measure of correlation between the results


in the oral and composition tests, justifying your choice of measure. Interpret the value you


obtain.


(b) Calculate the value of the most appropriate measure of correlation between the results in


the composition and literature tests, justifying your choice of measure. Interpret the value


you obtain.



1
Expert's answer
2022-03-20T06:40:05-0400

(a) The ranks of oral test and composition are:



Spearman's correlation coefficient:

"\bar{x}=5\\\bar{y}=5\\\sum{\left( x_i-\bar{x} \right) ^2}=58\\\sum{\left( y_i-\bar{y} \right) ^2}=60\\\sum{\left( x_i-\bar{x} \right) \left( y_i-\bar{y} \right)}=42\\\rho =\frac{\sum{\left( x_i-\bar{x} \right) \left( y_i-\bar{y} \right)}}{\sqrt{\sum{\left( x_i-\bar{x} \right) ^2}\sum{\left( y_i-\bar{y} \right) ^2}}}=\frac{42}{\sqrt{58\cdot 60}}=0.711967"

There is a strong positive correlation between the scores.

(b)



Pearson's correlation coefficient:

"\bar{x}=16\\\bar{y}=55.22222\\\sum{\left( x_i-\bar{x} \right) ^2}=136\\\sum{\left( y_i-\bar{y} \right) ^2}=1805.556\\\sum{\left( x_i-\bar{x} \right) \left( y_i-\bar{y} \right)}=473\\\rho =\frac{\sum{\left( x_i-\bar{x} \right) \left( y_i-\bar{y} \right)}}{\sqrt{\sum{\left( x_i-\bar{x} \right) ^2}\sum{\left( y_i-\bar{y} \right) ^2}}}=\frac{473}{\sqrt{136\cdot 1805.556}}=0.954523"

There is a strong positive correlation between the scores.